The respective elasticities of demand and supply for a product are -0.55 and 1. Further, it is observed that the price of this commodity is $80 whilst the prevailing quantity in the market is 950. By adopting a demand and supply concept of

Qd = a + bP
Qs = c + dP use the information provided to estimate the values of a,b,c and d.
1
Expert's answer
2020-09-24T11:05:19-0400

price elasticity of demand="\dfrac{\triangle Q^d}{\triangle P}*{P \over Q^d}"

"\Rightarrow -0.55=" "\dfrac{\triangle Q^d}{\triangle P}*{80 \over 950}"

"\Rightarrow" "\dfrac{\triangle Q^d}{\triangle P}" =-0.55*"{80 \over 950}"

=-6.53125

"Q^d= a+ bP"

"\Rightarrow" "\dfrac{\triangle Q^d}{\triangle P}" "=b"

"\therefore" "b=-6.53125"

"\Rightarrow Q^d = a+ (-6.53125)p"

taking p =80, Qd =950

"\Rightarrow 950=a+(-6.51325)(80)"

950 =a-522.5

a=950+522.5

a=1472.5

Qd =1472.5 -6.51325P


price elasticity of supply "=\dfrac{\triangle Q^s}{\triangle P}*{P \over Q^s}"

"\Rightarrow 1=\dfrac{\triangle Q^d}{\triangle P}*{80 \over 950}"

"\Rightarrow \dfrac{\triangle Q^s}{\triangle P}={950 \over 80}"

=11.875

"Q^s =c +dP"

"\Rightarrow" "\Rightarrow \dfrac{\triangle Q^s}{\triangle P} =d"

"\therefore" d=11.875

"Q^s =c +11.875p"

At p =80 , "Q^s =950"

950=c +11.875(80)

950=c+950

c=950-950

c=0

so Qs =0+11.875p

Qs =11.875p

"\therefore a=1472.5 , b=-6.53125 , c=0 ,d =11.875"